OpenStudy (anonymous):

Write the equation of the quadratic function with roots 0 and 2 and a vertex at (1, 5).Enter your answer as a function beginning with f(x). Do not write any spaces. Use ^ to type exponents. Leave your answer in the form:f(x)=(a)(x^2+bx+c) Sample answer:f(x)=(-1/8)(x^2+5x-12)

OpenStudy (anonymous):

since the roots are given at x=0 and x=2, the factored quadratic will look like: \(\large y=a(x-0)(x-2)=a(x^2-2x) \) \(\large y=a(x^2-2x) \) now you need to find "a".... since you know the vertex passes through (1, 5), plug this into that last equation and solve for a....

OpenStudy (anonymous):

so 1 is x and y is 5 so would it look like this 5=a(1^2-2(1) is that correct

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

so a = ????

OpenStudy (anonymous):

5

OpenStudy (anonymous):

good so now your can you get the equation in standard form y=ax^2 + bx + c ???

OpenStudy (anonymous):

i have no clue.. I got y=5 (1^2 then where does the b and c come in

OpenStudy (anonymous):

ohh... is this the way they want the answer: \(\large f(x)=(a)(x^2+bx+c) \) ????? or is it: \(\large f(x)=ax^2+bx+c \) ????

OpenStudy (anonymous):

the first one the f(x)=(a)(x^2+bx+c)

OpenStudy (anonymous):

the first one so: \(\large f(x)=(-5)(x^2-2x) \) is your answer.. double check that a=-5 not 5... is that right???

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

then ur done....